OpenStudy (anonymous):

The sides of a square that are 3 ft long increasingat a rate of 2 ft/min .how fast is the area that square increasing at that instant

OpenStudy (zzr0ck3r):

are you doing related rates in your class?

OpenStudy (anonymous):

This is the new function defined by the condition: \[f(x) = (2x+3)^{2}\] where x represents minutes elapsed. Expanding the equation: \[f(x) = 4x^{2}+12x+9\] Then the derivative would be: \[8x+12\] Mm. I'm not sure. Not an expert in calculus.. :s

OpenStudy (shubhamsrg):

let side be x now dx/dt = 2 (given) area = A =x^2 so dA/dt = d(x^2)/dt = 2x .(dx/dt) = 2x.2 since its asking you at that instant when x=3,,make that substitution and you'll get your ans..

OpenStudy (shubhamsrg):

@veximeer sorry buddy,,i'dd you're wrong..hmm..

OpenStudy (shubhamsrg):

i meant i'd guess**

OpenStudy (anonymous):

Me not an expert in calculus :))

OpenStudy (shubhamsrg):

nevermind ^_^